OpenStudy (anonymous):

solve for x -3|x+7|=-12

OpenStudy (anonymous):

-12/-3?

OpenStudy (shamim):

|x|=x or -x

OpenStudy (dan815):

|x+7|=-12/-3 |x+7|=4 an absolute value can be broken down into 2 equations like so, x+7 =4 or x+7=-4

OpenStudy (anonymous):

-3[x+7]=-12 x+7=4 and x+7=-4

OpenStudy (dan815):

since getting 4 or -4 in the absolute value will make it 4 after applying the absolute value function